Shown as follows is a segmented income statement for Drexel-Hall during the current month. 60 Sales Variable costs ContributiRequired A Required B Required c Total monthly sales for Drexel-Hall stores. expected in total monthly sales

Required A Required B Required C The monthly responsibility margin of Store Monthly Responsibility Margin Store 1 Store 2

Required A Required B Required C The companys monthly income from operations. Expected in monthly operating income
Answer #1

Answer: sales Store 1 = 600,000+ 54,000 654,000 stort 2 = 600,000 + 129,000 = 429,000 654,000+ 729,000 = 1,383,000 14:15 storStore s-lorea 101,730 monthly repubikathy wagen Responsbility marga : 80520 - Responsbility mergin (66,000) (be for elosing s